OpenStudy (anonymous):

Solve x2- 5x + 6 = 0 using the quadratic formula.

pooja195 (pooja195):

List your abc values

Nnesha (nnesha):

\[\huge\rm \frac{ - b \pm \sqrt{b^2 -4ac} }{ 2a }\] quadratic formula where abc values are \[\huge\rm Ax^2 +Bx+C=0\]

OpenStudy (anonymous):

?

OpenStudy (nuttyliaczar):

Imagine your equation is in the form Ax^2+Bx+C=0. What is your A, B, and C?

pooja195 (pooja195):

What are the A B C values

OpenStudy (nuttyliaczar):

Also sometimes you would have to manipulate your equation to get it in that form but in this case it's set up nicely for you

OpenStudy (anonymous):

a=1 b=5 and c=6

pooja195 (pooja195):

hmm a=1 b= -5 c=6

OpenStudy (anonymous):

cant i do x^2 +2x + 3x -5?

pooja195 (pooja195):

LEts stick to this its less complicated

pooja195 (pooja195):

\[\huge~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a } \]

pooja195 (pooja195):

fimiliar with this?

OpenStudy (anonymous):

x2- 5x + 6 = 0 x^2 - 2 + 3 + 6 (x^2-2x) (3x+6) divide by x and 2 so....

OpenStudy (anonymous):

yes

pooja195 (pooja195):

By factoring? Its asking to use the quadratic formula .

pooja195 (pooja195):

\[\huge~x=\frac{ -(-5) \pm \sqrt{(-5)^2-4(1)(6)} }{ 2(1) } \]

pooja195 (pooja195):

simplyfy this

OpenStudy (anonymous):

So the x2 can be written as 2x because they are communitive terms. Since the 2x and 5x are like terms, they can combine so 2x-5x = -3x. Ok. Now, since u want to use quadratic u need to assign an a b and c value. The a value is the coefficient of an x^2 in a quadratic equation. Since u don't have a quadratic equation, the coefficient is 0. So the a value is zero. The b value is the -3 and the c value is the number without an x (I don't remember what it was.) Plug the a, b and c values into the quadratic equation and solve. I hope that helps, because i am not good at explaining things. If u still don't understand, I'm sure Kahn academy has a video on it

OpenStudy (anonymous):

\[5\pm/2 \sqrt{-6}\]

OpenStudy (anonymous):

Oh. I got all the numbers wrong...

OpenStudy (anonymous):

Very sorry

pooja195 (pooja195):

\[\frac{ 5\pm \sqrt{1} }{ 2 }\]

pooja195 (pooja195):

simplfy

pooja195 (pooja195):

Set up 2 equations

OpenStudy (anonymous):

\[5\pm 1/2\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

The equation is 5+-(5^2 -4(1)(6))^1/2 /2(1). ^1/2 means square root

OpenStudy (anonymous):

jpro, can u stop talking

pooja195 (pooja195):

thats what i have above xD \[\frac{ 5+\sqrt{1} }{ 2}\] \[\frac{ 5-\sqrt{1} }{ 2}\] I did this because we have a plus and minus sign so we need 2 seperate equations

OpenStudy (anonymous):

so + that and - that

pooja195 (pooja195):

yea :)

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

x = 2 and x = 3 x = - 4 and x = - 3 x = 5 and x = - 3 x = - 2 and x = 4

OpenStudy (anonymous):

3 and 2

pooja195 (pooja195):

yep :)

pooja195 (pooja195):

Thats right good job :)

OpenStudy (anonymous):

i learned this along time ago the fist thing i learned, but what im doing now has nothing to with it...

pooja195 (pooja195):

lol what are you learning right now ?

OpenStudy (anonymous):

things like x^2 +6=180

pooja195 (pooja195):

Oh so like factoring?

pooja195 (pooja195):

Let me know if you need any help :)

pooja195 (pooja195):

\(\huge\color{orange}{Welcome~To~OpenStudy}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (anonymous):

not really were r kinda using a diffent methon with squareroots

pooja195 (pooja195):

oh right thats what its called :)